OpenStudy (rogue):

I know I have had this question before and got it right, but it's been a week or two since then and I can't remember or find anything in the notes I took. In the bullseye shown above, AB=BC=CD=DE, and

OpenStudy (rogue):

In the bullseye shown above, AB=BC=CD=DE, and AE=14. Calculate the difference in the areas of the blue and red rings. Round the answer to the nearest square inch.

OpenStudy (rogue):

I would greatly appreciate help!

OpenStudy (calculusxy):

It looks like as if the distances of each of the points, from A to B, B to C, C to D, and D to E, all looks like it's equal. There are also 4 intervals and from the most inner point (A) to the most outer point (E) is the difference of 14. Thus, I would say to divide 14 by 4 to figure out the distance of each interval, which would be 3.5. The formula for finding the area of a circle is \(\pi r^2\). The radius of the red ring is composed of two 3.5 distances (meaning 3.5 x 2). So it would be 7 as the radius of the red ring. So the area would be approximately 21.98 units^2. I am not sure whether you can leave it as that for the area of the red ring or you need to take out the portion of the yellow ring that covers the red ring. But I will do this just in case. The yellow ring's radius is 3.5. So 3.5 x 3.14 is approximately 10.99. So 21.98 - 10.99 = 10.99 would be the area of red ring without the without ring's portion. The blue ring would have the radius of 3.5^3 = 42.875. So 42.875 x 3.14 = 134.6275 which we can approximate it to 134.63 units^3. Then take out the portion covered by the red ring to get (134.63 - 21.98) 112.65 units^2. The difference would be 112.65 - 10.99 = 101.66 units^2. (I hope that this is correct. This is a new type of problem that I have tackled.)

OpenStudy (rogue):

THANK YOU SO MUCH

OpenStudy (calculusxy):

No problem :)